OpenStudy (zyberg):

Find integer triples of x, y and z (x; y; z) that are solutions to the system of equations: x^2 - y^2 - z^2 = 1; y + z - x = 3

OpenStudy (zyberg):

I tried to do traditional ways to solve this, but it is way too hard and I had achieved nothing, apart from finding very high powers of unknown numbers ;)

OpenStudy (bobo-i-bo):

You have 2 equations but three unknowns, therefore you have one degree of freedom meaning you get to arbitrarily assign a number to any variable... well almost :P Anyways, out of just pure randomness, let \(z=0\) and see whether you can find a solution from that. :)

OpenStudy (zyberg):

Solved it, going to post the answer after ~8 hours. It doesn't involve any "take z to be 0" (in fact, z isn't equal to 0 for integer solutions ;) ). Found this MSE question to be very useful: http://math.stackexchange.com/questions/1273844/how-to-solve-3-variable-in-2-equation

ganeshie8 (ganeshie8):

Nice, I'm looking forward to see the solution

OpenStudy (bobo-i-bo):

Oh right, sorry. I didn't read the bit about them having to be integer solutions. :P

OpenStudy (zyberg):

@ganeshie8, here's the solution: First we express x in terms of y and z, from the second equation: x = y + z - 3 Then we square x and get it to be equal to -> z^2 + y^2 + 2yz - 6z - 6y + 9 After that we just replace x in the first equation with that long sequence of numbers to get: z^2 + y^2 + 2yz - 6z - 6y + 9 - y^2 - z^2 = 1 As we can see, squares cancel out: 2yz - 6z - 6y + 8 = 0 Then, we find that we can factor this into: (y - 3)(2z - 6) = 10 10 = 5 * 2 or 10 * 1, so 2z-6 must equal either 1, 2, 5 or 10 (same for y-3). We notice that 2z - 6 = 2(z - 3), so it's always an even number, meaning that it can be either 2 or 10, leaving only 1 and 5 for y - 3. We apply this to find the solutions: If 2z - 6 = 10; z = 8, y = 4, x = 9; If 2z - 6 = 2; z = 4, y = 8, x = 9; However, we have overlooked the fact that 10 as well can be factored to negative numbers, so we have to check for 2z - 6 = -10 and 2z- 6 = -2: If 2z - 6 = -10; z = -2, y = 2, x = -3; If 2z - 6 = -2; z = 2, y = -2, x = -3; So, all the answers together: 1) x = 9, y = 4, z = 8; 2) x = 9, y = 8, z = 4; 3) x = -3, y = -2, z = 2; 4) x = -3, y = 2, z = -2.

ganeshie8 (ganeshie8):

Factorization did the magic! Clever !